What is the slope of a line perpendicular to a line with y = 3x - 5?
Answer by RAY100(1637)

y=3x-5
.
The slope intercept form is ,,,,,y=mx+b,,,
.
fitting the form,,,,m=3,,,,slope
.
to be perpendicular, a line must have negative reciprocal slope,
.
m(2) = -1/m(1)
.
in this case,,,m(2) =-1/3= (-1/3)
.